OpenStudy (anonymous):

I WILL FAN AND MEDAL, If f(x) = 2x + 3 and g(x) = x2 + 1,find f(g(3)). A)23 B)47 C)82 D)90

OpenStudy (anonymous):

@ajo21 @Ghostedly

OpenStudy (ghostedly):

A. 23.

OpenStudy (anonymous):

explain how

OpenStudy (ghostedly):

g(x) = x^2 + 1 g(3)= 3^2+1 = 10 f(x) = 2x + 3 f(7) = 2*10 + 3 =23

OpenStudy (anonymous):

okay thanks

OpenStudy (ghostedly):

Yup.